OpenStudy (anonymous):

Consider the sixth roots of unity and choose ALL correct responses A. The argument of any sixth root of unity is a multiple of 3 B. If z is a sixth root of unity, then z2 is also a sixth root of unity C. If z is a sixth root of unity, then z is also a 12th root of unity D. If z is a sixth root of unity, then z is a cube root of unity E. If z is a sixth root of unity, then −z is also a sixth root of unity

OpenStudy (anonymous):

Here are the six sixth root of unity: \[ \large \left\{1,e^{\frac{i \pi }{3}},e^{\frac{2 i \pi }{3}},-1,e^{-\frac{2 i \pi }{3}},e^{-\frac{i \pi }{3}}\right\} \] They are the solutions of \[ z^6 =1 \]

OpenStudy (anonymous):

It easy to check which statement is true.